Find the eccentricity of the ellipse given by: x^2/144 + y^2/63 = 1
Answer by lwsshak3(11628)

given: x^2/144 + y^2/63 = 1
Standard form of ellipse: x^2/a^2 + y^2/b^2 =1, a>b
a^2=144
a=12
b^2=63
b=sqrt(63)
c^2=a^2-b^2
c=sqrt(a^2-b^2)=sqrt(144-63)=sqrt(81)=9
eccentricity = c/a=9/12=.75
ans: eccentricity of given ellipse = .75